OpenStudy (anonymous):

write quotient in standard form 9+5i / 6-9i

OpenStudy (anonymous):

multiply both the top and the bottom by the conjugate of the denominator. The conjugate of a+bi = a-bi. Do that and see what happens to the denominator. :)

OpenStudy (anonymous):

btw, the standard form is a+bi

OpenStudy (anonymous):

by 6-9i ?

OpenStudy (anonymous):

6+9i. b in this case would be -9 so the conjugate is 6-(-9)i which is equal to 6+9i

OpenStudy (anonymous):

i mean 6+9i ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

yes , sorry

OpenStudy (anonymous):

whenever you want to get rid of the imaginary part of a complex denominator, multiply the top and bottom by the conjugate of the denominator.

OpenStudy (anonymous):

9+5i * 6+9i ? 6-9i * 6+9i ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I assume that the first line is the numerator and the second line is the denominator?

OpenStudy (anonymous):

Just making sure

OpenStudy (anonymous):

54 + 45i ? 36 + 81i ?

OpenStudy (anonymous):

yes first line is the numerator and the second line is the denominator

OpenStudy (anonymous):

Ok, you got the denominator right but (9+5i)*(6+9i) = 54 + 81i + 30i + 45

OpenStudy (anonymous):

wait, sorry. the denominator is wrong too. it is 36+81

OpenStudy (anonymous):

why not 81i ?

OpenStudy (anonymous):

i^2 = -1 because i = \[\sqrt{-1}\] so i^2 = \[\sqrt{-1}\] * \[\sqrt{-1}\] = -1

OpenStudy (anonymous):

agh sorry about not having it in one line

OpenStudy (anonymous):

its okay (: , im just trying to understand it

OpenStudy (anonymous):

do you understand how i^2 is -1?

OpenStudy (anonymous):

so i would have to add the top and it'll be 99 + 111i ?

OpenStudy (anonymous):

that it will result to the opposite ?

OpenStudy (anonymous):

\[(9+5i) / (6-9i) = [(9+5i)(6+9i)] / [(6-9i)(6+9i)]\] Then multiply it out. Remember that (a+b)(c+d) = ac + ad + bc + bd

OpenStudy (anonymous):

So \[(9+5i)(6+9i) = 9*6 + 9*9i + 6*5i + 5*9*i^2\]

OpenStudy (anonymous):

5*9*i^2 = -1*5*9. Just do the same thing for the denominator, too.

OpenStudy (anonymous):

so its not (9+5i)*(6+9i) = 54 + 81i + 30i + 45 ?

OpenStudy (anonymous):

suppose x+yi(x,y∈R),satisfies (9+5i)÷(6-9i)=a+bi , then (x+yi)*(6-9i)=9+5i. we must remember the logarithm of imaginaries as: (a+bi)(c+di)=(ac-bd)+(bc+ad)i. then 6x+9y=9, 6y-9x=5. so x=1/13, y=37/39. so 9+5i / 6-9i = 1/13 + 37/39 i

OpenStudy (anonymous):

That's right kathy. So in standard form it is (54+45) + (81+30)i = 99+111i Now try doing the denominator. Something interesting should happen.

OpenStudy (anonymous):

my method is a traditional one. Johnfreeman uses the idea of rationalization.

OpenStudy (anonymous):

yes thats the answer they give here as an option :) so to solve problems with i i would have to use logarithm of imaginaries?

OpenStudy (anonymous):

see this:

OpenStudy (anonymous):

to multiply aswell ?

OpenStudy (anonymous):

I think the easiest way to think of it at this point is to multiply the top and bottom by the conjugate because then you get a rational denominator.

OpenStudy (anonymous):

It should at least be good practice to try multiplying the denominator by its conjugate so you know what happens when you do that. It is probably how they want you to solve this anyway.

OpenStudy (anonymous):

clear, you can use tthe data in the problem to directly work it out: 9+5i / 6-9i=[9*6+5*(-9)]/[6^2+(-9)^2] + [5*6-9*(-9)]/[6^2+(-9)^2]*i = 1/13 + 37/39 i

OpenStudy (anonymous):

see, the answer's just the same.